WRITE THE EQUATION OF THE LINE THAT PASSES THRU THE GIVEN POINT AND IS PERPENDICULAR TO THE GIVEN LINE. WRITE THE ANSWER IN SLOPE INTERCEPT FORM.
(-4,2), y+2x = -17

Answer by Alan3354(69443)

WRITE THE EQUATION OF THE LINE THAT PASSES THRU THE GIVEN POINT AND IS PERPENDICULAR TO THE GIVEN LINE. WRITE THE ANSWER IN SLOPE INTERCEPT FORM.
(-4,2)
y+2x = -17
-----------------
y = -2x-17
The slope, m, is -2
The slope of lines perpendicular is the negative inverse, = 1/2
then use:
y-y1 = m*(x-x1) where (x1,y1) is the point
y-2 = (1/2)*(x+4)
y = (1/2)x + 4 (slope-intercept form)
x - 2y = -8 (standard form)